OpenStudy (anonymous):

2(3 – y) = 6 – 2(1 – 3y) y

OpenStudy (anonymous):

Distribute 2 into the parenthesis. 2 x 3 = ?

OpenStudy (andrewthecookie):

Disturbite

OpenStudy (anonymous):

6

OpenStudy (andrewthecookie):

2 x y =?

OpenStudy (anonymous):

Yes, and 2 x y = 2y. So the left side is 6 - 2y.

OpenStudy (anonymous):

Next, distribute 4 into the other parenthesis. Can you do this yourself? @pootypie

OpenStudy (andrewthecookie):

Now disturbite on the other side

OpenStudy (anonymous):

4y? And Yeah

OpenStudy (anonymous):

So multiply 2 x 1 and 2 x 3y.

OpenStudy (anonymous):

2*1=2 And 2*3=6

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Yes, so the equation is now 6 – 2y = 6 – 2 – 6y

OpenStudy (anonymous):

U like the saints @AnswerMyQuestions

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

@pootypie Subtract 6 - 2.

OpenStudy (anonymous):

i did it's 2

OpenStudy (anonymous):

6 - 2...

OpenStudy (anonymous):

I Was Writing It On Paper

OpenStudy (anonymous):

OMG I love them to !

OpenStudy (anonymous):

@Mesopretty Just Tag Him In A Status